OpenStudy (anonymous):

Write in lowest terms: 5 3/8 / 2 3/4 A. 10 9/32 B. 3 1/2 C. 2 1/2 D. 1 21/22

OpenStudy (anonymous):

@perl @Awolflover1 @kidrah69

OpenStudy (texaschic101):

(5 3/8) / (2 3/4) --- turn to improper fractions (43/8) / (11/4) = 43/8 * 4/11 = 172/88 = 1 84/88 = 1 21/22